OpenStudy (amenah8):

find dy/dx of y=sqrt(lnx) and of y=cos(ln2x^2)

OpenStudy (misty1212):

hI!!

OpenStudy (misty1212):

chain rule for both

OpenStudy (amenah8):

i ended up with (1)/(x^1/2) but i don't think it's right

OpenStudy (misty1212):

\[(f\circ g)'=f(g)\times g'\] with \[f(x)=\sqrt{x}, f'(x)=\frac{1}{2\sqrt{x}}, g(x)=\ln(x), g'(x)=\frac{1}{x}\]

OpenStudy (misty1212):

no that is not right

OpenStudy (misty1212):

lets cut to the chase \[\frac{d}{dx}\sqrt{f(x)} =\frac{f'(x)}{2\sqrt{f(x)}}\]

OpenStudy (amenah8):

why is it 2\[\sqrt{f(x)}\]

OpenStudy (misty1212):

the derivative of the square root of x is one over two square root of x

OpenStudy (misty1212):

therefore, by the chain rule, the derivative of the square root of something, is one over two square root of something, times the derivative of something

OpenStudy (amenah8):

so what if f prime of x? x^.5?

OpenStudy (misty1212):

for example, the derivative of \[f(x)=\sqrt{\sin(x)}\] is \[\frac{\cos(x)}{2\sqrt{\sin(x)}}\]

OpenStudy (ac3):

same thing .5=1/2

OpenStudy (ac3):

x^(1/2) = sqrt(x)

OpenStudy (misty1212):

forget exponential notation, just memorize it

OpenStudy (ac3):

here let me show you why.

OpenStudy (misty1212):

\[f(x)=\sqrt{x}\\ f'(x)=\frac{1}{2\sqrt{x}}\] it is always the same, like \(7\times 8=56\) alway s

OpenStudy (misty1212):

the square root is a very very (very) common function, just know it so you can access it any time forget the power rule, that is just a gimmick

OpenStudy (amenah8):

so the derivative of y = sqrt(lnx) is: 1/2sqrt(x) ?

OpenStudy (misty1212):

hmm no

OpenStudy (amenah8):

(1/\[\sqrt{x}\])

OpenStudy (amenah8):



OpenStudy (ac3):